# Upstream history quantification and scale-decomposed energy analysis for weak-to-strong adverse-pressure-gradient turbulent boundary layers

### v1

@@ 4 Summary and conclusions
In addition to fortifying our inferences on the effects of accumulated PG on TBL statistics, Case 2 offers insight into the role of cumulative disequilibration, which leads to a delayed turbulence response. The rapid increase in the APG strength (${\rm d}{\beta_{C}}/{\rm d}{Re_{\theta}}$) in DNS22 manifests into a delayed turbulence response resulting in statistics with lower energisation than expected for the local $\beta$ value, particularly affecting the outer region of the TBL. Case 2 thus highlights the limitation of the accumulated PG parameter in capturing history effects in situations involving rapid disequilibration, which induces a pronounced delayed response in turbulence.

### v2

@@ 4 Summary and conclusions
In addition to fortifying our inferences on the effects of accumulated PG on TBL statistics, Case 2 offers insight into the role of local disequilibration, which leads to a delayed turbulence response. The rapid increase in the APG strength (${\rm d}{\beta_{C}}/{\rm d}{Re_{\theta}}$) in DNS22 manifests into a delayed turbulence response resulting in statistics with lower energisation than expected for the local $\beta$ value, particularly affecting the outer region of the TBL. Case 2 thus highlights the limitation of the accumulated PG parameter in capturing history effects in situations involving rapid disequilibration, which induces a pronounced delayed response in turbulence. Case DISEQ further highlights the impact of local disequilibration on turbulent scales, revealing that the near-wall small-scale region adapts rapidly to the changing pressure gradient, while other regions exhibit a delayed response.